OpenStudy (anonymous):

Find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function. f(x)=-2x^2+2x+9 (I need the x-coordinate and the y-coordinate and the equation for the line of symmetry.) Please? and thank you. :)

hero (hero):

You need all of that, huh?

OpenStudy (saifoo.khan):

hmm.

OpenStudy (anonymous):

According to the question in my homework...

hero (hero):

And you can't find one of those by yourself?

OpenStudy (saifoo.khan):

it's a Combo question.

OpenStudy (anonymous):

Yes, saif knows. :)

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

I tried using the wolframalpha site, but I can only come up with the graph. I don't know what the other properties translate to.

OpenStudy (saifoo.khan):

vertex = (1/2 , 19/2) sym = x=1/2 max, 19/2

OpenStudy (anonymous):

oooooooooooooooooook! woo! Lifesaver! :D

hero (hero):

What would you do without Saifoo?

OpenStudy (saifoo.khan):

Thanks for the lifesaver thing. =)

OpenStudy (anonymous):

Walk to the campus that is 5 miles away without a car and try to find a study group and leave my infant daughter at home with my cancer patient father... Why?

OpenStudy (anonymous):

He has been a bigger help than he will ever know. :)

OpenStudy (anonymous):

= no car, no family

hero (hero):

What's a snot rocket?

OpenStudy (anonymous):

You really wanna know? lol

hero (hero):

I'm scared :P

OpenStudy (anonymous):

I have a different sense of humor and sometimes it's gross. hahaha! ;p

OpenStudy (shadowfiend):

Hero, you need to reconsider the tone in which you answer questions like this. We want members of our community to feel welcome, not talked down to. Feel free to guide them, not chastise them.